80% of the people who purchase ramen from Angeline are women


Answer the following in percent format (4 decimal places each).

If 9 of these customers are randomly selected, then the probability that all of them are women is Blank 1%.


If 10 of these customers are randomly selected, then the probability that there are at least 7 women is Blank 2%


1
Expert's answer
2021-05-07T12:37:10-0400

"p = 0.8 \\

q = 1 -p = 1 -0.8 = 0.2"

If 9 of these customers are randomly selected, then the probability that all of them are women is

"n = 9 \\

P(X=x) = \frac{n!}{x!(n-x)!} \times p^x \times (q)^{n-x} \\

P(X=9) = \frac{9!}{9!(9-9)!} \times 0.8^9 \times (0.2)^{9-9} \\

= 1 \times 0.134217 \times 1 \\

= 0.134217 \\

= 13.4217 \; \%"

If 10 of these customers are randomly selected, then the probability that there are at least 7 women is

"n = 10 \\

P(X≥7) = P(X=7) + P(X=8) + P(X=9) + P(X=10) \\

P(X=7) = \frac{10!}{7!(10-7)!} \times 0.8^7 \times (0.2)^{10-7} \\

= 120 \times 0.2097 \times 0.008 \\

= 0.201312 \\

P(X=8) = \frac{10!}{8!(10-8)!} \times 0.8^8 \times (0.2)^{10-8} \\

= 45 \times 0.1677 \times 0.04 \\

= 0.30186 \\

P(X=9) = \frac{10!}{9!(10-9)!} \times 0.8^9 \times (0.2)^{10-9} \\

= 10 \times 0.134217 \times 0.2 \\

= 0.268434 \\

P(X=10) = \frac{10!}{10!(10-10)!} \times 0.8^10 \times (0.2)^{10-10} \\

= 1 \times 0.1073741 \times 1 \\

= 0.1073741 \\

P(X≥7) = 0.201312 + 0.30186+ 0.268434 + 0.1073741 = 0.878980 \\

= 87.8980 \; \%"